If ski lift raises 100 passengers averaging 660 N in weight, a machine raises passengers averaging 667 N in weight to a height of 150 m in 60.0 s, at constant speed what aveage power is required of the force making the lift
Answer #1

Hi,

Hope you are doing well.

First of all let's find out the total mass in the lift which is given by,

11 = 100 × aug

Where,

aug is the average force per person (N).

g is the acceleration due to gravity (m/s2).

Given that,

Average force per person, F_{avg}=667\;N

Acceleration due to gravity, g=9.8\;m/s^{2}

\therefore Total mass in the lift is given by,

M=100\times\frac{F_{avg}}{g}=100\times\frac{667\;N}{9.8\;m/s^{2}}

\mathbf{\therefore M=6806.1224499\;kg}

We know that, When the lift reaches a particular height, it will gain a potential energy which is given by,

U=Mgh

Where,

M is the Total mass in the lift (kg).

g is the acceleration due to gravity (m/s2).

h is the height reached by the lift (m).

Given that,

Total mass in the lift, M=6806.1224499\;kg

Acceleration due to gravity, g=9.8\;m/s^{2}

Height reached by the lift, h=150\;m

\therefore Potential energy gained by the lift is given by,

U=Mgh=6806.1224499\;kg\times9.8\;m/s^{2}\times150\;m

\mathbf{\therefore U=10,005,000\;J}

We know that, according to work energy theorem, the work done by the lift to lift is converted to potential energy.

\therefore Work done by the lift is given by,

W=U=10,005,000\;J

We know that, Power required by the lift is given by,

P=W\times t

Where,

t is the time taken (s).

Given that,

Time consumed for the lifting, t=60.0\;s

\therefore Power required by the lift is given by,

P=W\times t=10,005,000\;J\times 60.0\;s

\mathbf{\therefore P=600.3\;MW}

\left (\because 10^{6}\;W=1\;MW \right )

Hope this helped for your studies. Keep learning. Have a good day.

Feel free to clear any doubts at the comment section.


Please don't forget to give a thumbs up.

Thank you. :)